# A Note on Restricted Selection Set from Random Interval

### v1

@@ 2.7. Event probabilities: unrestricted and mean-restricted ranges
We now turn to the range of probabilities $\mathbf{P}(y^{*}\in A)$ as $y^{*}$ varies over selections of $Y$, with and without a mean restriction. Let $A$ be a fixed Borel subset of $\mathbb{R}$.
@@
Both bounds are sharp in the sense that there exist measurable selections whose probabilities are arbitrarily close to the respective endpoints.
@@
This is a measurable random closed set with nonempty values, so by measurable selection there exists $y^{+}$ with $y^{+}(\omega)\in G_{+}(\omega)$ almost surely (see [CastaingValadier1977, Molchanov2005]). Then $y^{+}(\omega)\in A$ whenever $Y(\omega)\cap A\neq\emptyset$, and necessarily $y^{+}(\omega)\notin A$ when $Y(\omega)\cap A=\emptyset$. Hence
@@
The next result uses a Lagrangian and an interchange of $\sup$ and expectation in the spirit of [Artstein1974, Molchanov2005]; convex duality arguments are standard, see e.g. [Rockafellar1970].
@@
Moreover, $U$ is concave and nonincreasing in $\kappa$, while $L$ is convex and nondecreasing.
@@
We sketch the argument for $U(\kappa)$; the case of $L(\kappa)$ is analogous.
For each $\lambda\in\mathbb{R}$, consider the Lagrangian functional
@@
For fixed $\lambda$, maximizing $\mathcal{L}(y,\lambda)$ over $y\in\operatorname{\mathbf{Sel}}(Y)$ under the pointwise constraint $y(\omega)\in Y(\omega)$ is equivalent to maximizing
@@
over all selections $y$ of $Y$. By an interchange theorem for random sets (e.g. Artstein [Artstein1974], see also [Molchanov2005]),
@@
with the understanding that the second term is -$\infty$ if $Y(\omega)\cap A=\emptyset$. For $\lambda<0$ one similarly obtains $\max\{\lambda y_{L}(\omega),1+\lambda a^{-}(\omega)\}$. Thus
@@
On the other hand, for each $y\in\operatorname{\mathbf{Sel}}(Y\mid\kappa)$, $\mathcal{L}(y,\lambda)=\mathbf{P}(y\in A)$ for all $\lambda$, so
@@
By convex duality (the feasible set $\{\mathbb{E}[y]:y\in\operatorname{\mathbf{Sel}}(Y)\}$ is the compact interval $\mathbb{E}y_{L},\mathbb{E}y_{U}]$ and the objective is linear in $y$),
see, e.g., Rockafellar [Rockafellar1970]. Combining this with the calculation above yields the expression for $U(\kappa)$ in (2.16). The expression for $L(\kappa)$ is obtained by applying the same reasoning to the minimization problem for $\mathbf{P}(y\in A)$ subject to $\mathbb{E}[y]=\kappa$. The concavity/convexity and monotonicity properties follow directly from the dual envelopes. ∎
@@
For each $\lambda\in\mathbb{R}$, let $y_{\lambda}(\omega)\in\arg\max_{x\in Y(\omega)}\big(\mathbf{1}\{x\in A\}+\lambda x\big)$ with measurable tie-breaking on the boundary. Then
$y_{\lambda}(\omega)=\begin{cases}y_{U}(\omega),&\lambda>0,\ Y(\omega)\cap A=\emptyset,\\[4.0pt] y_{U}(\omega)-\min\{\Delta^{+}(\omega),\,1/\lambda\},&\lambda>0,\ Y(\omega)\cap A\neq\emptyset,\\[4.0pt] y_{L}(\omega)+\min\{\Delta^{-}(\omega),\,1/(-\lambda)\},&\lambda<0,\\[4.0pt] \text{any element of }Y(\omega),&\lambda=0.\end{cases}$ (2.17)
The map $\lambda\mapsto\mathbf{E}[y_{\lambda}]$ is continuous and nondecreasing, with $\lim_{\lambda\to-\infty}\mathbf{E}[y_{\lambda}]=\mathbf{E}y_{L}$ and $\lim_{\lambda\to+\infty}\mathbf{E}[y_{\lambda}]=\mathbf{E}y_{U}$. Hence for each $\kappa\in[\mathbf{E}y_{L},\mathbf{E}y_{U}]$ there exists $\lambda^{\star}$ such that $\mathbf{E}[y_{\lambda^{\star}}]=\kappa$, and any such $y_{\lambda^{\star}}$ attains $U(\kappa)$.
@@
Fix $\lambda>0$. On each $\omega$, the function $x\mapsto\mathbf{1}_{A}(x)+\lambda x$ is affine with slope $\lambda>0$ outside $A$, and affine with the same slope plus a unit jump on $A$. If $Y(\omega)\cap A=\emptyset$, the maximizer is clearly $y_{U}(\omega)$, giving the first line of (2.17). If $Y(\omega)\cap A\neq\emptyset$, denote
$a^{+}(\omega)=\sup(Y(\omega)\cap A),\qquad\Delta^{+}(\omega)=y_{U}(\omega)-a^{+}(\omega)\geq 0.$
On $Y(\omega)\cap A$, the maximum of $1+\lambda x$ is attained at $x=a^{+}(\omega)$, giving value $1+\lambda a^{+}(\omega)$. On $Y(\omega)$, the maximum of $\lambda x$ is $\lambda y_{U}(\omega)$. Thus the maximizer is $y_{U}(\omega)$ if $\lambda y_{U}(\omega)\geq 1+\lambda a^{+}(\omega)$, i.e. $\lambda\Delta^{+}(\omega)\geq 1$; otherwise it is $a^{+}(\omega)$. This yields
@@
$y_{\lambda}(\omega)=\begin{cases}y_{U}(\omega),&\lambda\Delta^{+}(\omega)\geq 1,\\ a^{+}(\omega),&\lambda\Delta^{+}(\omega)<1,\end{cases}$
which is equivalent, for $Y(\omega)\cap A\neq\emptyset$, to
$y_{\lambda}(\omega)=y_{U}(\omega)-\min\{\Delta^{+}(\omega),\,1/\lambda\}$
when $\lambda>0$. The case $\lambda<0$ is analogous, with roles of $y_{U}$ and $y_{L}$ reversed and $a^{-}(\omega)=\inf(Y(\omega)\cap A)$.
As $\lambda\to+\infty$, $1/\lambda\to 0$ and hence $\min\{\Delta^{+},1/\lambda\}\to 0$ almost surely on $\{Y\cap A\neq\emptyset\}$, while $y_{\lambda}(\omega)=y_{U}(\omega)$ when $Y(\omega)\cap A=\emptyset$, so $y_{\lambda}\to y_{U}$ and $\mathbb{E}[y_{\lambda}]\to\mathbb{E}[y_{U}]$. As $\lambda\to$-$\infty$, $y_{\lambda}\to y_{L}$ and $\mathbb{E}[y_{\lambda}]\to\mathbb{E}[y_{L}]$. Monotonicity and continuity of $\lambda\mapsto\mathbb{E}[y_{\lambda}]$ follow from the monotone dependence of the thresholds on $\lambda$ and dominated convergence. Existence of a $\lambda^{\star}$ with $\mathbb{E}[y_{\lambda^{\star}}]=\kappa$ for any $\kappa\in[\mathbb{E}y_{L},\mathbb{E}y_{U}]$ then follows by the intermediate value theorem.
Finally, for such a $\lambda^{\star}$, the dual expression in Theorem 2.13 yields
@@
$U(\kappa)\geq\mathbf{P}(y_{\lambda^{\star}}\in A)=\mathbb{E}\big[\mathbf{1}_{A}(y_{\lambda^{\star}})\big],$
while the Lagrangian argument shows that equality holds at the calibrated $\lambda^{\star}$. ∎

### v2

@@ 2.7. Event probabilities: unrestricted and mean-restricted ranges
We now turn to the range of probabilities $\mathbf{P}(y\in A)$ as $y$ varies over selections of $Y$, with and without a mean restriction. Let $A$ be a fixed Borel subset of $\mathbb{R}$.
@@
Both bounds are attained by measurable selections.
@@
This is a measurable random closed set with nonempty values, so by measurable selection there exists $y^{+}$ with $y^{+}(\omega)\in G_{+}(\omega)$ almost surely (see [3, 7]). Then $y^{+}(\omega)\in A$ whenever $Y(\omega)\cap A\neq\emptyset$, and necessarily $y^{+}(\omega)\notin A$ when $Y(\omega)\cap A=\emptyset$. Hence
@@
The next result uses a Lagrangian and an interchange of $\sup$ and expectation in the spirit of [1, 7]; convex duality arguments are standard, see e.g. [8].
@@
Moreover, $U$ is concave in $\kappa$, while $L$ is convex.
@@
We prove the formula for $U(\kappa)$; the argument for $L(\kappa)$ is analogous.
For each $\lambda\in\mathbb{R}$, consider the Lagrangian
@@
For fixed $\lambda$, maximizing $\mathcal{L}(y,\lambda)$ over $y\in\operatorname{\mathbf{Sel}}(Y)$ is equivalent to maximizing
@@
subject to the pointwise constraint $y(\omega)\in Y(\omega)$. By the standard interchange theorem for random closed sets (see, e.g., [1, 7]),
@@
with the convention that the second term is -$\infty$ when $Y(\omega)\cap A=\varnothing$. If $\lambda<0$, similarly,
$\sup_{x\in Y(\omega)}f_{\lambda}(\omega,x)=\max\{\lambda y_{L}(\omega),\,1+\lambda a^{-}(\omega)\}.$
Hence
@@
On the other hand, if $y\in\operatorname{\mathbf{Sel}}(Y\mid\kappa)$, then $\mathbb{E}[y]=\kappa$, so
$\mathcal{L}(y,\lambda)=\mathbf{P}(y\in A)\qquad\text{for every }\lambda\in\mathbb{R}.$
Therefore
@@
It remains to justify the minimax interchange. Define the feasible set
$\mathcal{Y}:=\bigl\{(\mathbb{E}[y],\,\mathbb{E}[\mathbf{1}_{A}(y)]):\,y\in\operatorname{\mathbf{Sel}}(Y)\bigr\}\subset\mathbb{R}^{2}.$
Then
$U(\kappa)=\sup\{\,p:(\kappa,p)\in\mathcal{Y}\,\}=\sup_{(m,p)\in\mathcal{Y}}\inf_{\lambda\in\mathbb{R}}\bigl[p+\lambda(m-\kappa)\bigr].$
We first show that $\mathcal{Y}$ is convex. Let
$(m_{i},p_{i})\in\mathcal{Y},\qquad i=1,2,$
so there exist selections $y_{i}\in\operatorname{\mathbf{Sel}}(Y)$ with
$m_{i}=\mathbb{E}[y_{i}],\qquad p_{i}=\mathbb{E}[\mathbf{1}_{A}(y_{i})].$
Fix $\theta\in[0,1]$. Pass to the atomless product extension
where $\lambda$ is Lebesgue measure, and let $U$ be the coordinate uniform random variable on $[0,1]$, independent of all random objects on $\Omega$. Extend $Y,y_{1},y_{2}$ to $\bar{\Omega}$ by ignoring the second coordinate, and define
$\bar{y}(\omega,u):=\mathbf{1}_{\{u\leq\theta\}}\,y_{1}(\omega)+\mathbf{1}_{\{u>\theta\}}\,y_{2}(\omega).$
Then $\bar{y}(\omega,u)\in Y(\omega)$ for all $(\omega,u)$, so $\bar{y}\in\operatorname{\mathbf{Sel}}(Y)$ on the product space. By independence of $U$,
$\mathbb{E}_{\bar{\mathbf{P}}}[\bar{y}]=\theta\,\mathbb{E}[y_{1}]+(1-\theta)\,\mathbb{E}[y_{2}]=\theta m_{1}+(1-\theta)m_{2},$
and
$\mathbb{E}_{\bar{\mathbf{P}}}[\mathbf{1}_{A}(\bar{y})]=\theta\,\mathbb{E}[\mathbf{1}_{A}(y_{1})]+(1-\theta)\,\mathbb{E}[\mathbf{1}_{A}(y_{2})]=\theta p_{1}+(1-\theta)p_{2}.$
Thus
$\theta(m_{1},p_{1})+(1-\theta)(m_{2},p_{2})\in\mathcal{Y},$
after identifying the product extension with the original atomless setting. Hence $\mathcal{Y}$ is convex.
Next, $\mathcal{Y}$ is bounded, because every selection satisfies $y_{L}\leq y\leq y_{U}$ a.s., so
$\mathbb{E}[y]\in[\mathbb{E}y_{L},\mathbb{E}y_{U}],\qquad\mathbb{E}[\mathbf{1}_{A}(y)]\in[0,1].$
Thus
$\mathcal{Y}\subset[\mathbb{E}y_{L},\mathbb{E}y_{U}]\times[0,1].$
We now note that the upper boundary function
$\phi(m):=\sup\{\,p:(m,p)\in\mathcal{Y}\,\},\qquad m\in[\mathbb{E}y_{L},\mathbb{E}y_{U}],$
is concave, since $\mathcal{Y}$ is convex. In particular,
$U(\kappa)=\phi(\kappa).$
A basic fact from finite-dimensional convex analysis is that every finite concave function on an interval is the pointwise infimum of its affine majorants. Equivalently,
$\phi(\kappa)=\inf_{\lambda\in\mathbb{R}}\sup_{(m,p)\in\mathcal{Y}}\bigl[p+\lambda(m-\kappa)\bigr].$
Therefore
$U(\kappa)=\inf_{\lambda\in\mathbb{R}}\sup_{(m,p)\in\mathcal{Y}}\bigl[p+\lambda(m-\kappa)\bigr].$
Since each $(m,p)\in\mathcal{Y}$ is of the form
$(m,p)=\bigl(\mathbb{E}[y],\,\mathbb{E}[\mathbf{1}_{A}(y)]\bigr)\quad\text{for some }y\in\operatorname{\mathbf{Sel}}(Y),$
this becomes
$U(\kappa)=\inf_{\lambda\in\mathbb{R}}\sup_{y\in\operatorname{\mathbf{Sel}}(Y)}\Bigl\{\mathbb{E}[\mathbf{1}_{A}(y)]+\lambda(\mathbb{E}[y]-\kappa)\Bigr\}=\inf_{\lambda\in\mathbb{R}}\Bigl\{\mathbb{E}[\Psi(\lambda,\cdot)]-\lambda\kappa\Bigr\}.$
This proves the first identity in (2.17).
The formula for $L(\kappa)$ is obtained in the same way by replacing the supremum problem for $\mathbf{P}(y\in A)$ with the corresponding infimum problem. Finally, since $U$ is the pointwise infimum of affine functions of $\kappa$, it is concave, and since $L$ is the pointwise supremum of affine functions of $\kappa$, it is convex. ∎
@@
Assume $A$ is closed. For each $\lambda\in\mathbb{R}$, define the argmax multifunction
$Y_{\lambda}(\omega):=\arg\max_{x\in Y(\omega)}\bigl(\mathbf{1}_{\{x\in A\}}+\lambda x\bigr).$
Then $Y_{\lambda}$ has nonempty closed values and admits measurable selections. Moreover, its graph is measurable.
If $\lambda>0$, then
$Y_{\lambda}(\omega)=\begin{cases}\{y_{U}(\omega)\},&Y(\omega)\cap A=\emptyset,\\[4.0pt] \{y_{U}(\omega)\},&Y(\omega)\cap A\neq\emptyset,\ \lambda\Delta^{+}(\omega)>1,\\[4.0pt] \{a^{+}(\omega)\},&Y(\omega)\cap A\neq\emptyset,\ \lambda\Delta^{+}(\omega)<1,\\[4.0pt] \{a^{+}(\omega),\,y_{U}(\omega)\},&Y(\omega)\cap A\neq\emptyset,\ \lambda\Delta^{+}(\omega)=1.\end{cases}$
Hence, outside the tie set
$T_{\lambda}^{+}:=\{\omega:Y(\omega)\cap A\neq\emptyset,\ \lambda\Delta^{+}(\omega)=1\},$
every measurable maximizer satisfies
$y_{\lambda}(\omega)=\begin{cases}y_{U}(\omega),&Y(\omega)\cap A=\emptyset,\\[4.0pt] y_{U}(\omega)-\min\{\Delta^{+}(\omega),\,1/\lambda\},&Y(\omega)\cap A\neq\emptyset.\end{cases}$ (2.18)
Note that on $\{Y\cap A\neq\emptyset\}\setminus T_{\lambda}^{+}$, the formula evaluates to $a^{+}(\omega)$ when $\lambda\Delta^{+}(\omega)<1$ and to $y_{U}(\omega)$ when $\lambda\Delta^{+}(\omega)>1$ (since $\min\{\Delta^{+},1/\lambda\}=\Delta^{+}=y_{U}-a^{+}$ in the first case and $1/\lambda$ in the second, with $1/\lambda<\Delta^{+}\leq y_{U}-y_{L}$). When $Y(\omega)\cap A=\emptyset$, the unique maximizer is $y_{U}(\omega)$ as stated. If $\lambda<0$, then
$Y_{\lambda}(\omega)=\begin{cases}\{y_{L}(\omega)\},&Y(\omega)\cap A=\emptyset,\\[4.0pt] \{y_{L}(\omega)\},&Y(\omega)\cap A\neq\emptyset,\ (-\lambda)\Delta^{-}(\omega)>1,\\[4.0pt] \{a^{-}(\omega)\},&Y(\omega)\cap A\neq\emptyset,\ (-\lambda)\Delta^{-}(\omega)<1,\\[4.0pt] \{a^{-}(\omega),\,y_{L}(\omega)\},&Y(\omega)\cap A\neq\emptyset,\ (-\lambda)\Delta^{-}(\omega)=1.\end{cases}$
Hence, outside the tie set
$T_{\lambda}^{-}:=\{\omega:Y(\omega)\cap A\neq\emptyset,\ (-\lambda)\Delta^{-}(\omega)=1\},$
every measurable maximizer satisfies
$y_{\lambda}(\omega)=\begin{cases}y_{L}(\omega),&Y(\omega)\cap A=\emptyset,\\[4.0pt] y_{L}(\omega)+\min\{\Delta^{-}(\omega),\,1/(-\lambda)\},&Y(\omega)\cap A\neq\emptyset.\end{cases}$ (2.19)
If $\lambda=0$, then
$Y_{0}(\omega)=\begin{cases}Y(\omega)\cap A,&Y(\omega)\cap A\neq\emptyset,\\ Y(\omega),&Y(\omega)\cap A=\emptyset.\end{cases}$
In particular, $Y_{0}$ is again a measurable closed-valued multifunction.
Moreover, for every $\kappa\in[\mathbb{E}y_{L},\mathbb{E}y_{U}]$, there exist $\lambda^{\star}\in\mathbb{R}$ and $y^{\star}\in\operatorname{\mathbf{Sel}}(Y_{\lambda^{\star}})$ such that
$\mathbb{E}[y^{\star}]=\kappa,\qquad\mathbf{P}(y^{\star}\in A)=U(\kappa).$
@@
Fix $\lambda\in\mathbb{R}$ and define
$f_{\lambda}(\omega,x):=\mathbf{1}_{A}(x)+\lambda x,\qquad x\in Y(\omega).$
We first identify the argmax multifunction $Y_{\lambda}$ pointwise.
Case 1: $\lambda>0$. If $Y(\omega)\cap A=\emptyset$, then $f_{\lambda}(\omega,x)=\lambda x$ on $Y(\omega)$, hence the unique maximizer is $y_{U}(\omega)$.
Now assume $Y(\omega)\cap A\neq\emptyset$. Since $A$ is closed and $Y(\omega)$ is compact,
$a^{+}(\omega):=\sup\bigl(Y(\omega)\cap A\bigr)\in Y(\omega)\cap A.$
On $Y(\omega)\cap A$, the map $x\mapsto 1+\lambda x$ is increasing, so its maximum is attained at $a^{+}(\omega)$ and equals $1+\lambda a^{+}(\omega)$. On $Y(\omega)\setminus A$, the map $x\mapsto\lambda x$ is increasing, so its maximum is attained at $y_{U}(\omega)$ and equals $\lambda y_{U}(\omega)$. Thus the maximizers are determined by comparing
$\lambda y_{U}(\omega)\quad\text{and}\quad 1+\lambda a^{+}(\omega).$
Since
$\lambda y_{U}(\omega)\geq 1+\lambda a^{+}(\omega)\iff\lambda\bigl(y_{U}(\omega)-a^{+}(\omega)\bigr)\geq 1\iff\lambda\Delta^{+}(\omega)\geq 1,$
we obtain exactly the stated description of $Y_{\lambda}(\omega)$ for $\lambda>0$. Outside the tie set $T_{\lambda}^{+}$, this is equivalent to (2.18).
Case 2: $\lambda<0$. If $Y(\omega)\cap A=\emptyset$, then $f_{\lambda}(\omega,x)=\lambda x$ on $Y(\omega)$, and since $\lambda<0$ the unique maximizer is $y_{L}(\omega)$.
Now assume $Y(\omega)\cap A\neq\emptyset$. Since $A$ is closed and $Y(\omega)$ is compact,
$a^{-}(\omega):=\inf\bigl(Y(\omega)\cap A\bigr)\in Y(\omega)\cap A.$
On $Y(\omega)\cap A$, the map $x\mapsto 1+\lambda x$ is decreasing, so its maximum is attained at $a^{-}(\omega)$ and equals $1+\lambda a^{-}(\omega)$. On $Y(\omega)\setminus A$, the map $x\mapsto\lambda x$ is decreasing, so its maximum is attained at $y_{L}(\omega)$ and equals $\lambda y_{L}(\omega)$. Hence the maximizers are determined by
$\lambda y_{L}(\omega)\geq 1+\lambda a^{-}(\omega)\iff(-\lambda)\bigl(a^{-}(\omega)-y_{L}(\omega)\bigr)\geq 1\iff(-\lambda)\Delta^{-}(\omega)\geq 1.$
This gives the stated description of $Y_{\lambda}(\omega)$ for $\lambda<0$. Outside the tie set $T_{\lambda}^{-}$, this is equivalent to (2.19).
Case 3: $\lambda=0$. Then $f_{0}(\omega,x)=\mathbf{1}_{A}(x)$, so the maximizers are exactly
$Y_{0}(\omega)=\begin{cases}Y(\omega)\cap A,&Y(\omega)\cap A\neq\emptyset,\\ Y(\omega),&Y(\omega)\cap A=\emptyset.\end{cases}$
In all three cases, $Y_{\lambda}(\omega)$ is a nonempty closed subset of the compact interval $Y(\omega)$. The preceding pointwise characterization shows that $Y_{\lambda}$ is built from the measurable objects $y_{L}$, $y_{U}$, $^{+}, aa^{-}$, $\Delta^{+}$, and $\Delta^{-}$ by measurable case distinctions; hence $\operatorname{graph}(Y_{\lambda})\in\mathcal{A}\otimes\mathcal{B}(\mathbb{R})$. Therefore the Kuratowski–Ryll-Nardzewski theorem yields $\operatorname{\mathbf{Sel}}(Y_{\lambda})\neq\emptyset$.
Now let $y_{\lambda}\in\operatorname{\mathbf{Sel}}(Y_{\lambda})$. By construction,
$f_{\lambda}(\omega,y_{\lambda}(\omega))=\sup_{x\in Y(\omega)}f_{\lambda}(\omega,x)=\Psi(\lambda,\omega)\qquad\text{a.s.}$
Therefore
$\mathbb{E}\big[\mathbf{1}_{A}(y_{\lambda})+\lambda y_{\lambda}\big]=\mathbb{E}[\Psi(\lambda,\cdot)].$ (2.20)
In particular, every measurable selection of $Y_{\lambda}$ attains the pointwise-maximized Lagrangian.
Fix $\kappa\in[\mathbb{E}y_{L},\mathbb{E}y_{U}]$ and define
$g(\lambda):=\mathbb{E}[\Psi(\lambda,\cdot)]-\lambda\kappa.$
By Theorem 2.13,
$U(\kappa)=\inf_{\lambda\in\mathbb{R}}g(\lambda).$
Since for each $\omega$ the map $\lambda\mapsto\Psi(\lambda,\omega)$ is the supremum of affine functions of $\lambda$, $g$ is convex and lower semicontinuous.
If $\kappa=\mathbb{E}y_{U}$, then $y_{U}\in\operatorname{\mathbf{Sel}}(Y\mid\kappa)$ and
$U(\kappa)=\mathbf{P}(y_{U}\in A),$
so $y^{\star}:=y_{U}$ is optimal. Likewise, if $\kappa=\mathbb{E}y_{L}$, then $y^{\star}:=y_{L}$ is optimal. Hence it remains to treat the interior case
$\kappa\in(\mathbb{E}y_{L},\mathbb{E}y_{U}).$
For $\lambda\geq 0$,
$\Psi(\lambda,\omega)\geq\lambda y_{U}(\omega),$
so
$g(\lambda)\geq\lambda(\mathbb{E}y_{U}-\kappa)\xrightarrow[\lambda\to+\infty$$\infty$.
For $\lambda<0$,
$\Psi(\lambda,\omega)\geq\lambda y_{L}(\omega),$
so
$g(\lambda)\geq\lambda(\mathbb{E}y_{L}-\kappa)\xrightarrow[\lambda\to-\infty$$\infty$.
Thus $g$ is coercive and attains its minimum at some $\lambda^{\star}\in\mathbb{R}$.
We next show that one can calibrate the mean at $\lambda^{\star}$.
For $\lambda>0$, define the lower and upper tie-breaking selections
@@
$y_{\lambda}^{-}(\omega):=\begin{cases}a^{+}(\omega),&\omega\in T_{\lambda}^{+},\\ \text{the unique maximizer in }Y_{\lambda}(\omega),&\omega\notin T_{\lambda}^{+},\end{cases}$
and
$y_{\lambda}^{+}(\omega):=\begin{cases}y_{U}(\omega),&\omega\in T_{\lambda}^{+},\\ \text{the unique maximizer in }Y_{\lambda}(\omega),&\omega\notin T_{\lambda}^{+}.\end{cases}$
For $\lambda<0$, define analogously
$y_{\lambda}^{-}(\omega):=\begin{cases}y_{L}(\omega),&\omega\in T_{\lambda}^{-},\\ \text{the unique maximizer in }Y_{\lambda}(\omega),&\omega\notin T_{\lambda}^{-},\end{cases}$
and
$y_{\lambda}^{+}(\omega):=\begin{cases}a^{-}(\omega),&\omega\in T_{\lambda}^{-},\\ \text{the unique maximizer in }Y_{\lambda}(\omega),&\omega\notin T_{\lambda}^{-}.\end{cases}$
Finally, for $\lambda=0$, choose measurable selections $y_{0}^{-},y_{0}^{+}\in\operatorname{\mathbf{Sel}}(Y_{0})$ such that
$y_{0}^{-}(\omega)=\begin{cases}\inf(Y(\omega)\cap A),&Y(\omega)\cap A\neq\emptyset,\\ y_{L}(\omega),&Y(\omega)\cap A=\emptyset,\end{cases}$
and
$y_{0}^{+}(\omega)=\begin{cases}\sup(Y(\omega)\cap A),&Y(\omega)\cap A\neq\emptyset,\\ y_{U}(\omega),&Y(\omega)\cap A=\emptyset.\end{cases}$
Set
$m_{-}(\lambda):=\mathbb{E}[y_{\lambda}^{-}],\qquad m_{+}(\lambda):=\mathbb{E}[y_{\lambda}^{+}].$
By construction,
$m_{-}(\lambda)\leq\mathbb{E}[y]\leq m_{+}(\lambda)\qquad\text{for every }y\in\operatorname{\mathbf{Sel}}(Y_{\lambda}).$
We claim that
$g^{\prime}_{-}(\lambda)=m_{-}(\lambda)-\kappa,\qquad g^{\prime}_{+}(\lambda)=m_{+}(\lambda)-\kappa.$ (2.21)
Indeed, for each fixed $\omega$, the one-sided derivatives of $\lambda\mapsto\Psi(\lambda,\omega)$ are exactly the smallest and largest slopes of the active affine pieces, namely
$D_{-}\Psi(\lambda,\omega)=y_{\lambda}^{-}(\omega),\qquad D_{+}\Psi(\lambda,\omega)=y_{\lambda}^{+}(\omega).$
Moreover,
$|D_{\pm}\Psi(\lambda,\omega)|\leq|y_{L}(\omega)|+|y_{U}(\omega)|\qquad\text{a.s.},$
and the right-hand side is integrable. Hence dominated convergence yields
$\frac{d^{-}}{d\lambda}\mathbb{E}[\Psi(\lambda,\cdot)]=\mathbb{E}[D_{-}\Psi(\lambda,\cdot)]=m_{-}(\lambda),$
and
$\frac{d^{+}}{d\lambda}\mathbb{E}[\Psi(\lambda,\cdot)]=\mathbb{E}[D_{+}\Psi(\lambda,\cdot)]=m_{+}(\lambda).$
Subtracting the derivative of $\lambda\kappa$ gives (2.21).
Since $\lambda^{\star}$ minimizes the convex function $g$, we have
$g^{\prime}_{-}(\lambda^{\star})\leq 0\leq g^{\prime}_{+}(\lambda^{\star}).$
Using (2.21),
$m_{-}(\lambda^{\star})\leq\kappa\leq m_{+}(\lambda^{\star}).$
@@
We now construct a measurable $y^{\star}\in\operatorname{\mathbf{Sel}}(Y_{\lambda^{\star}})$ with $\mathbb{E}[y^{\star}]=\kappa$.
If $\lambda^{\star}>0$, then every maximizer agrees outside $T_{\lambda^{\star}}^{+}$, while on $T_{\lambda^{\star}}^{+}$ the only possible values are $a^{+}$ and $y_{U}$. On this tie set,
$y_{U}-a^{+}=\Delta^{+}=1/\lambda^{\star}\qquad\text{a.s. on }T_{\lambda^{\star}}^{+}.$
Hence
$m_{+}(\lambda^{\star})-m_{-}(\lambda^{\star})=\frac{1}{\lambda^{\star}}\mathbf{P}(T_{\lambda^{\star}}^{+}).$
If $\mathbf{P}(T_{\lambda^{\star}}^{+})=0$, then $m_{-}(\lambda^{\star})=m_{+}(\lambda^{\star})$ and necessarily $\kappa=m_{-}(\lambda^{\star})$; take $y^{\star}:=y_{\lambda^{\star}}^{-}$. If $\mathbf{P}(T_{\lambda^{\star}}^{+})>0$, choose
$\theta:=\frac{\kappa-m_{-}(\lambda^{\star})}{m_{+}(\lambda^{\star})-m_{-}(\lambda^{\star})}\in[0,1].$
By non-atomicity there exists a measurable set $B\subseteq T_{\lambda^{\star}}^{+}$ with
$\mathbf{P}(B)=\theta\,\mathbf{P}(T_{\lambda^{\star}}^{+}).$
Define
$y^{\star}:=\mathbf{1}_{B}\,y_{\lambda^{\star}}^{+}+\mathbf{1}_{B^{c}}\,y_{\lambda^{\star}}^{-}.$
Then $y^{\star}\in\operatorname{\mathbf{Sel}}(Y_{\lambda^{\star}})$ and
$\mathbb{E}[y^{\star}]=m_{-}(\lambda^{\star})+\frac{1}{\lambda^{\star}}\mathbf{P}(B)=m_{-}(\lambda^{\star})+\theta\bigl(m_{+}(\lambda^{\star})-m_{-}(\lambda^{\star})\bigr)=\kappa.$
If $\lambda^{\star}<0$, the same argument applies on $T_{\lambda^{\star}}^{-}$, where
$a^{-}-y_{L}=\Delta^{-}=1/(-\lambda^{\star})\qquad\text{a.s. on }T_{\lambda^{\star}}^{-}.$
Thus one again obtains a measurable $y^{\star}\in\operatorname{\mathbf{Sel}}(Y_{\lambda^{\star}})$ with $\mathbb{E}[y^{\star}]=\kappa$.
If $\lambda^{\star}=0$, then
$Y_{0}(\omega)=\begin{cases}Y(\omega)\cap A,&Y(\omega)\cap A\neq\emptyset,\\ Y(\omega),&Y(\omega)\cap A=\emptyset,\end{cases}$
is a one-dimensional random closed set. Since
$m_{-}(0)\leq\kappa\leq m_{+}(0),$
the one-dimensional Aumann expectation formula yields a selection $y^{\star}\in\operatorname{\mathbf{Sel}}(Y_{0})$ with $\mathbb{E}[y^{\star}]=\kappa$.
Finally, since $y^{\star}\in\operatorname{\mathbf{Sel}}(Y_{\lambda^{\star}})$, it is a measurable pointwise maximizer. Hence by (2.20),
$\mathbb{E}\big[\mathbf{1}_{A}(y^{\star})+\lambda^{\star}y^{\star}\big]=\mathbb{E}[\Psi(\lambda^{\star},\cdot)].$
Using $\mathbb{E}[y^{\star}]=\kappa$, we get
$\mathbf{P}(y^{\star}\in A)=\mathbb{E}[\Psi(\lambda^{\star},\cdot)]-\lambda^{\star}\kappa=g(\lambda^{\star})=\inf_{\lambda\in\mathbb{R}}g(\lambda)=U(\kappa),$
where the last equality is Theorem 2.13. Thus $y^{\star}$ attains $U(\kappa)$. ∎